# Accelerated first detection in discrete-time quantum walks using sharp restarts

### v1

@@ Abstract
We divide the search for an optimal $\tau$ into two cases: (1) when the parity of the target position is known, and (2) when it is unknown. Case (1) can be simplified to a search for even $\delta$s in which both even and odd measurement periods can compete. In contrast, for case (2), we must only use odd $\tau$ values to ensure detection for all $\delta$s. Fig.3(b) and (c) show the minimum expected FDPT as a function of $\delta$s for cases (1) and (2), respectively. The plots reveal that for even target positions, $\tau=4$ yields the smallest values of minimum $\langle\text{FDPT}\rangle_{r}$ fiercely competing with $\tau=1$ and $2$ whose plots, as expected, are identical. All values of $\tau>2$ (except $\tau=4$) fall far behind in this race of target search. Hence, when the parity of the target position is known, we can turn it into an even value by shifting the origin if required. Then $\tau=4$ is the optimal measurement period that enables the fastest search. In case (2), where we search for a target of unknown parity, Fig.3(c) reveals that $\tau=1$ is optimal. We now turn to the reasons behind these observations.
@@
To explain these results, we examine an alternative expression for $\langle\text{FDPT}\rangle_{r}$ relevant to our analysis:
@@
$\langle\text{FDPT}\rangle_{r}=\tau\langle n_{f}\rangle=\tau[\sum_{\eta=1}^{\infty}\eta F_{(\eta\bmod r)}],$ (4)
@@
where $F_{0}\equiv F_{r}$ and $\eta$ denotes $\eta-th$ measurement step, with $n=\tau\eta$. Here, $F_{\eta}=(1-p_{1})...(1-p_{\eta-1})p_{\eta}$, as defined earlier. The first-order term with the highest contribution in the expansion of $F_{\eta}$ is $p_{\eta}$, followed by the second-order term, $p_{\eta}\sum_{i=1}^{\eta-1}p_{i}$, and so forth. However, approximating $F_{\eta}$ as $p_{\eta}$ is sufficient to explain the observed results.
@@
Why do walks with high $\tau$ values perform poorly compared to those with lower $\tau$s? This is because high values of $\tau$ allow the quantum walker to spread farther across the lattice between measurements [$n=\tau\eta$]. In high $\tau$ walks, compared to those with smaller $\tau$, the finite, conserved total probability is distributed over a larger domain, causing $p_{\eta}\sim F_{\eta}$ to take smaller values at low $\eta$, even with restarts at the ORT. Eq.4 indicates that walks with low $F_{\eta}$ at small $\eta$s yield higher $\langle\text{FDPT}\rangle_{r}$ than those with larger $F_{\eta}$ at similar $\eta$ values.
@@
However, if higher $\tau$s, which spread probability more widely across the lattice, generally increase $\langle\text{FDPT}\rangle_{r}$, then how does $\tau=4$ outperforms $\tau=1,2,$ and $3$ for even $\delta$ values? To answer this, two key points are relevant: (1) In general DTQWs, including those in this study, the probability of finding the walker at the position $x_{end}^{\pm}$ is very small. Here, $x_{end}^{\pm}$ represents the farthest lattice point with a non-zero chance of finding the walker; (2) Monitored-DTQW with $\tau=1$ or $2$ and $\delta>0$ ($\delta<0$) has $x_{end}^{+}\leq\delta$ ($x_{end}^{-}\geq\delta$). Alternatively, assuming $\delta>0$, the first time $p(x=\delta)$ takes a non-zero value, projective measurement |$\ket{\delta}\bra{\delta}$ | with a negative output changes the walker’s quantum state such that $x_{end}$ turns from $\delta$ to $\delta-2$. Thereafter, with only $1$ or $2$ steps between measurements ($\tau=1$ or $2$), the walker can evolve only as far as $x_{end}^{+}=\delta$, where it again faces the same outcome.
@@
Points (1) and (2) explain why, for even $\delta$ values, $\tau=4$ outperforms $\tau=1$ and $2$, despite the broader probability spread. With $\tau=4$, the walker moves beyond $\delta$, allowing its $p_{\eta}$ to exceed the limited $p_{\eta}=p(x_{end}^{\pm})$ achievable by $\tau=1$ or $2$. Contrastingly, walks with $\tau=3$ lose to those with $\tau=1,2$, and $4$, even after allowing its walker to go past $\delta$. This is due to redundant alternate measurements that make no change in the walker’s state, effectively allowing 6-step evolution and resulting in a broader probability spread that counteracts the benefits of moving past $\delta$.
@@
Conclusions — In this work, we explored the role of sharp-restarts, initial states, and measurement periods in expediting the quantum hitting times of DTQWs. First, we highlight the key findings: Monitored DTQWs with sharp restart significantly outperform their continuous-time quantum counterpart and optimally restarted continuous-time and discrete-time classical walks. The coin degree of freedom can be leveraged to tailor DTQWs with faster hitting times when the target’s position is known ($\delta\geq 0$ or $\delta\leq 0$). Notably, the mean FDPTs of restarted DTQWs exhibit oscillatory behavior, a hallmark of quantum behavior absent in classical walks. For higher measurement periods, the optimum time to restart the walk does not increase in a strictly monotonic fashion as the target moves farther from the origin. Furthermore, knowledge of the target’s parity can be utilized to get an even more efficient search by selecting measurement period $\tau=4$. The combined knowledge of the target’s side and parity enables highly efficient searches. Even without such prior information, initializing the walk with a symmetric coin state and setting the measurement period to one still ensures better performance than classical and continuous-time quantum walks. With the ability to control and engineer periodic Kumar et al. (2018) and aperiodic DTQW Lo Gullo et al. (2017), one can foresee an important role of sharp restart in modeling and accelerating quantum dynamics in complex quantum systems and quantum networks with directed targets.

### v2

@@ III Results
In DTQWs, the probability of finding the walker at all even (odd) positions is 0 at odd (even) steps of the walk [Eq. 7]. While for odd $\tau$s (say $\tau=3$), we measure the walker’s position at both odd and even steps alternatively (at $\eta=3,6,9,...$), for even $\tau$s, we can measure position only at even steps. Thus, for even $\tau$s, detection at odd $\delta$ is impossible. These observations not only suggest that setting an odd $\tau$ is a safer option when target’s parity is unknown, but also that if it is known, setting an even $\tau$ could be more beneficial. This is because when $\delta$’s parity is known, it can always be turned even by shifting the origin. Then, for even $\tau$s, we have finite detection probability at every measurement step, whereas for odd $\tau$s, we definitely miss detection at every alternate odd measurement step. This will be demonstrated later, with some nuances. However, first, let us move from $P_{det}(n)$ to expected FDPT under restart, denoted by $\langle\text{FDPT}\rangle_{r}$, which is a more robust measure of the efficiency of target detection.
Let $n_{f}$ be the number of measurements, and $\mathcal{R}$ be the number of restarts until first hitting. If we restart, as before, after every $r$th measurement step, $\langle n_{f}\rangle=r\langle\mathcal{R}\rangle+\langle\tilde{n}\rangle$, where $0\leq\mathcal{R}<\infty$ and $1\leq\tilde{n}\leq r$. From here onward, we set the discrete time step $\varepsilon=1$ (without loss of generality), thereby simplifying the relation to $\langle\text{FDPT}\rangle_{r}=\Delta T\langle n_{f}\rangle=\tau\langle n_{f}\rangle$ where [55, 42, 12]
$\begin{split}\langle n_{f}\rangle&=\sum_{\mathcal{R}=0}^{\infty}\sum_{\tilde{n}=1}^{r}(r\mathcal{R}+\tilde{n})[(1-P_{det}(r))^{\mathcal{R}}F_{\tilde{n}}]\\ &=\frac{1-P_{det}(r)}{P_{det}(r)}r+\sum_{\tilde{n}=1}^{r}\frac{(\tilde{n})F_{\tilde{n}}}{P_{det}(r)}.\end{split}$ (8)
The final equality follows from derivations for CTQW [39].
In Fig. 2(b), we plot $\langle\text{FDPT}\rangle_{r}$ as a function of restart time $t_{r}=r\Delta T=r\tau(\because\varepsilon=1)$, keeping $\delta=10$ fixed for all walks. We see an oscillatory behavior in the mean FDPTs explicit to quantum walks. The results also support the claim that DTQWs initialized with a spin state |$\ket{s}$ that skews the PPD toward the side opposite the target outperform those with other initial states. Specifically, for $\tau=1$, the average first detection passage time, $\langle\text{FDPT}\rangle_{r}$, is lowest when |$\ket{s}=\ket{1}$ and highest when |$\ket{s}=\ket{0}$. This occurs because the target site $\delta=10$ is on the right (positive) side of the origin, and initializing with |$\ket{s}=\ket{1}$ produces a left-skewed PPD—i.e., the peak of the PPD is closer to the target. Consequently, the probability of detecting the walker at $\delta=10$ is higher for DTQWs initialized with |$\ket{s}=\ket{1}$ compared to those with |$\ket{s}=\ket{0}$, where the PPD peaks away from the target. The increased detection probability for |$\ket{s}=\ket{1}$ reduces the FDPT on average, as reflected in Fig. 2(b). When |$\ket{s}$ is fixed at $\ket{+}$, $\tau=4$ yields the lowest expected FDPT under restart, followed closely by $\tau=1$ and $2$, with $\tau=3$ performing worst among them. We will compare the efficiency of different $\tau$s in detail later.
Now, we note another interesting result: for the same $\delta$ ($10$) and |$\ket{s}$ ($\ket{+}$), $\tau=2$ closely bounds the $\langle\text{FDPT}\rangle_{r}$ plot of $\tau=1$ from below. More generally, for all even $\delta$s, $\tau=2\tau^{\prime}$ tightly bounds $\tau^{\prime}$ whenever $\tau^{\prime}$ is odd. This is expected when we compare the evolution of the DTQW for an odd $\tau^{\prime}$ with its immediate even successor $2\tau^{\prime}$. We find that for odd $\tau^{\prime}$, where we measure walker’s position at $x=\delta$ at times $n\Delta T=\tau^{\prime},2\tau^{\prime},3\tau^{\prime},$ and so on. Since, $\delta$ is set even (here, $\delta=10$), measurements at the alternate odd steps ($\tau^{\prime},3\tau^{\prime},5\tau^{\prime},...)$ do not alter the walker’s quantum state, making both walks essentially the same. The minor difference arises due to their PPD following the equation: $p_{2\tau^{\prime}}(x=\delta|n)=p_{\tau^{\prime}}(x=\delta|2n)$. Here, $p_{2\tau^{\prime}}$ and $p_{\tau^{\prime}}$ is the probability of detecting the walker at $\delta$ during the $n$th and $2n$th measurement (i.e., effectively at same times $t_{1}=n\Delta T_{1}=n(2\tau^{\prime})=2n(\tau^{\prime})=2n\Delta T_{2}=t_{2}$) for DTQWs with measurement periods $2\tau^{\prime}$ and $\tau^{\prime}$, respectively. We conclude that, if the target location’s parity is known, selecting an even $2\tau^{\prime}$ is preferable, as using the odd half $\tau^{\prime}$ results in redundant measurements at every alternate step.
Before finding the most efficient $\tau$, let us examine how the expected FDPT under restart varies with $\delta$. In Fig. 2(c), keeping constant $\tau=1$, we plot $\langle\text{FDPT}\rangle_{r}$ as a function of restart time $t_{r}=r\Delta T=r\tau$ for different $\delta$s. In both Fig. 2 (b) and (c), the y-coordinate of the red points indicate minimum $\langle\text{FDPT}\rangle_{r}$, while their x-coordinate represent the optimum restart time. The optimum restart time (ORT), i.e., the restart time that minimizes $\langle\text{FDPT}\rangle_{r}$, increases as $\delta$ grows. One expects this result because to reach $\delta$ far from the origin, the walker needs to evolve for a longer time, necessitating a larger restart time. However, this increase in ORT with $\delta$ may not be strictly monotonic.
To illustrate this, we plot ORT as a function of even $\delta$s in Fig. 3(a). We keep only even $\delta$s to allow even $\tau$s along with the odd ones. For odd $\delta$s and even $\tau$s, detection is impossible, causing ORT to diverge to infinity. The plot shows that for $\tau=1$, ORT increases monotonically with $\delta$, as also observed in Fig. 2(c). The $\tau=2$ plot lies precisely on the top of $\tau=1$, further demonstrating the $\tau^{\prime}-2\tau^{\prime}$ equivalence for even $\delta$s. As $\tau$ increases from $3$ to $5$, sharper peaks and valleys emerge in the plots. However, the average slopes of all the $\tau$ values are similar, indicating that as $\delta$ increases, ORT generally increases as well. Finally, we examine how varying $\tau$ affects the search for different target positions to identify the optimal measurement period.
[…]
@@
We divide the search for an optimal $\tau$ into two cases: (1) when the parity of the target position is known, and (2) when it is unknown. Case (1) can be simplified to a search for even $\delta$s in which both even and odd measurement periods can compete. In contrast, for case (2), we must only use odd $\tau$ values to ensure detection for all $\delta$s. Fig. 3(b) and (c) show the minimum expected FDPT as a function of $\delta$s for cases (1) and (2), respectively. The plots reveal that for even target positions, $\tau=4$ yields the smallest values of minimum $\langle\text{FDPT}\rangle_{r}$ fiercely competing with $\tau=1$ and $2$ whose plots, as expected, are identical. All values of $\tau>2$ (except $\tau=4$) fall far behind in this race of target search. Hence, when the parity of the target position is known, we can turn it into an even value by shifting the origin if required. Then $\tau=4$ is the optimal measurement period that gives the smallest values of minimum $\langle\text{FDPT}\rangle_{r}$, enabling the fastest search. In case (2), where we search for a target of unknown parity, Fig. 3(c) reveals that $\tau=1$ is optimal. We now turn to the reasons behind these observations.
@@
To explain these results, we examine the expression for $\langle\text{FDPT}\rangle_{r}$ as a function of $\Delta T$. Keeping $\varepsilon=1$ as before such that $\Delta T=\tau$, we have
@@
$\begin{split}\langle\text{FDPT}\rangle_{r}(\tau)&=\tau\langle n_{f}\rangle\\ &=\frac{1-P_{det}(r)}{P_{det}(r)}(r\tau)+\sum_{\tilde{n}=1}^{r}\frac{(\tilde{n})F_{\tilde{n}}}{P_{det}(r)}\tau.\end{split}$ (9)
@@
Here, $F_{\tilde{n}}=(1-p_{1})...(1-p_{\tilde{n}-1})p_{\tilde{n}}$, as defined earlier. The first-order term with the highest contribution in the expansion of $F_{n}$ is $p_{n}$, followed by the second-order term, $p_{n}\sum_{i=1}^{n-1}p_{n}$, and so forth. However, approximating $F_{n}$ as $p_{n}$ is sufficient to explain the observed results. Similarly, the leading contribution in the expansion of $P_{det}(n)$ is $\sum_{i=1}^{n}p_{i}$, thus we approximate $P_{det}(n)\sim\sum_{i=1}^{n}p_{i}$.
@@
Why do walks with high $\tau$ values perform poorly compared to those with lower $\tau$s? In other words, why walks with higher $\tau$ exhibit higher minimum $\langle\text{FDPT}\rangle_{r}$ values (see Fig. 3(b) and (c))? This occurs because a larger $\tau$ allows the quantum walker to spread further across the lattice between measurements. For example, after 2 null measurements, a DTQW walker with $\tau=5$ reaches $x=2\times 5=10$, which is five times farther than one with $\tau=1$, reaching $x=2$. Hence, in high $\tau$ walks, compared to those with smaller $\tau$, the finite, conserved total probability is distributed over a larger domain, causing $p_{n}$ to take smaller values even with restarts. Now, keeping $\tau$ constant, if we minimize Eq. 9 with respect to $r$, we obtain the minimum $\langle\text{FDPT}\rangle_{r}$ for a given $\tau$. The restart time $r\tau$ after minimization becomes the ORT, $r^{*}\tau$, which remains nearly the same for walks with different $\tau$ (see Fig. 3(a)). Moreover, setting $r=r^{*}$, and noting that $F_{n}\sim p_{n}$ and $P_{det}(n)\sim\sum_{i=1}^{n}p_{i}$, a straightforward analysis shows that both terms in Eq. 9 increase with $\tau$, implying that the minimum $\langle\text{FDPT}\rangle_{r}$ rises as we increase $\tau$. In summary, walks with higher $\tau$ perform poorly compared to those with lower $\tau$.
@@
However, if higher $\tau$s, which spread probability more widely across the lattice, generally increase $\langle\text{FDPT}\rangle_{r}$, then how does $\tau=4$ outperforms $\tau=1,2,$ and $3$ for even $\delta$ values? To answer this, two key points are relevant: (1) In general DTQWs, including those in this study, the probability of finding the walker at the position $x_{end}^{\pm}$ is very small. Here, $x_{end}^{\pm}$ represents the farthest lattice point with a non-zero chance of finding the walker; (2) Monitored-DTQW with $\tau=1$ or $2$ and $\delta>0$ ($\delta<0$) has $x_{end}^{+}\leq\delta$ ($x_{end}^{-}\geq\delta$). Alternatively, assuming $\delta>0$, the first time probability of measuring the walker at $x=\delta$ becomes non-zero, a null measurement changes the walker’s quantum state such that $x_{end}$ turns from $\delta$ to $\delta-2$. Thereafter, with only $1$ or $2$ steps between measurements ($\tau=1$ or $2$), the walker can evolve only as far as $x_{end}^{+}=\delta$, where it again faces the same outcome.
@@
Points (1) and (2) explain why, for even $\delta$ values, $\tau=4$ outperforms $\tau=1$ and $2$, despite the broader probability spread. With $\tau=4$, the walker moves beyond $\delta$, allowing its $p_{n}$ to exceed the limited $p_{n}=p(x_{end}^{\pm})$ achievable by $\tau=1$ or $2$. Contrastingly, walks with $\tau=3$ lose to those with $\tau=1,2$, and $4$, even after allowing its walker to go past $\delta$. This is due to redundant alternate measurements that make no change in the walker’s state, effectively allowing 6-step evolution and resulting in a broader probability spread that counteracts the benefits of moving past $\delta$.
We note that the results here are obtained for the coin $\hat{\textsf{C}}(\theta=\pi/4)$ [Eq. 1]; however, they fundamentally depend on the nature of PPDs produced by the DTQWs. Consequently, with appropriate modifications, these results are generic and remain valid even if a different coin (e.g., the Hadamard coin) is used, provided that the analysis accounts for the corresponding PPDs.
IV Conclusion
@@
In this work, we explored the role of sharp-restarts, initial states, and measurement periods in expediting the quantum hitting times of DTQWs. First, we highlight the key findings: Monitored DTQWs with sharp restart significantly outperform their continuous-time quantum counterpart and optimally restarted continuous-time and discrete-time classical walks. As shown in restarted CTQWs [39], restart helps in evading dark states in the DTQWs context as well. The coin degree of freedom can be leveraged to tailor DTQWs with faster hitting times when the target’s position relative to the origin is known (i.e., $\delta\geq 0$ or $\delta\leq 0$). Notably, the mean FDPTs of restarted DTQWs exhibit oscillatory behavior, a hallmark of quantum behavior absent in classical walks. For higher measurement periods, the optimum time to restart the walk does not increase in a strictly monotonic fashion as the target moves farther from the origin. Furthermore, knowledge of the target’s parity can be utilized to get an even more efficient search by selecting measurement period $\tau=4$. The combined knowledge of the target’s side and parity enables highly efficient searches. Even without such prior information, initializing the walk with a symmetric coin state and setting the measurement period to one still ensures better performance than classical and continuous-time quantum walks.
With the ability to control and engineer periodic [56] and aperiodic DTQW [57], one can foresee an important role of sharp restart in modeling and accelerating quantum dynamics in complex quantum systems and quantum networks with directed targets.